OpenStudy (anonymous):

two friends are renting an apartment. they pay the landlord the first and last month's rent. The landlord also also requires them to pay an additional half month's rent for a security deposit. . the total amount due is 1600. what is the monthly rent. linear equations

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@peachpi @Mehek14

OpenStudy (anonymous):

@izuru

Vocaloid (vocaloid):

let x = one month's rent x + x + 0.5x = 1600 solve for x

OpenStudy (anonymous):

0.7x

Vocaloid (vocaloid):

not quite... remember that x = 1x x + x + 0.5x = (1+1+0.5)x = ?

OpenStudy (anonymous):

2.5x?

Vocaloid (vocaloid):

right. so, 2.5x = 1600 what is x?

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

ohh its 640

Vocaloid (vocaloid):

right

Vocaloid (vocaloid):

so $640 is our answer

OpenStudy (anonymous):

thanks. :)